The charge of 1 electron is 1.602x10^-19 C, coulombs is the derived SI unit of electrical charge. Electrical current is measured in units of ampere, the fundamental SI unit of electrical current, 1 ampere which corresponds to 1 C s^-1 of electrons passing a stationary point. Calculate the number of electrons per second corresponding to 1 ampere of electrical current that is 1 coulomb.(C)..
1
Expert's answer
2018-07-17T01:52:01-0400
Solution:
When 1 Coulomb charge flow through a wire in 1 second then the current through the wire is 1 ampere.
I = Q/t
1 Ampere = 1 Coulomb /1 Second
Charge on 1 electron = 1.602∙10^-19 Coulom
By unitary method,
If 1.602∙10-19 Coulomb / Second (Ampere) = Current by 1 electron
then, 1(Coulomb / Second) or (Ampere) = 1 / (1.602∙10^-19) electrons
i.e, N = 6.25∙10^18 electrons.